how much paper is needed to make a cylindrical paper cup with a base diameter of 10 cm and height of 12 cm
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

radius = 10/2 = 5 cm
slant height = sqrt(5^2 + 12^2) = 13 cm

surface area = pi * r * sqrt(r^2 + h^2)
= pi * 5 * 13
= 65pi

(the circular base is not counted because that part isn't covered by paper)

Answer by ikleyn(52803)

.
how much paper is needed to make a cylindrical paper cup with a base diameter of 10 cm and height of 12 cm
~~~~~~~~~~~~~~


To tutor @Penguin


The problem asks about CYLINDRICAL cups --- not about conical cups . . .